the value of k such that P(X<k) = 0.2578


1
Expert's answer
2021-09-23T00:16:52-0400

Full problem:


Given the normally distributed variable X with mean 18 and standard deviation 2.5, find

(a) P(X<15):

(h) P(17< X<21);

(c) the value of k such that P(X<k) = 0.2578;

(d) the value of k such that P(X> A) = 0.1539.


"\mu=18 \\

\sigma=2.5"

(a)

"P(X<15) = P(Z< \frac{15-18}{2.5}) \\

=P(Z< -1.2) \\

= 0.1150"

(b)

"P(17<X<21) = P(X<21) -P(X<17) \\

=P(Z< \frac{21-18}{2.5}) -P(Z< \frac{17-18}{2.5}) \\

=P(Z< 0.8571) -P(Z< -0.4) \\

= 0.8042 -0.3445 \\

= 0.4597"

(c)

"P(X<k) = 0.2578 \\

P(Z< \frac{k-18}{2.5}) =0.2578 \\

\frac{k-18}{2.5} = -0.65 \\

k-18 = 2.5 \times (-0.65) \\

k=18 -1.625 \\

k=16.375"

(d)

"P(X>k) = 0.1539 \\

P(X<k) = 1 -0.1539 = 0.8461 \\

P(Z< \frac{k-18}{2.5}) =0.8461 \\

\frac{k-18}{2.5} = 1.02 \\

k-18=2.5 \times 1.02 \\

k = 18 +2.55 \\

k=20.55"